# ViLCo-Bench: VIdeo Language COntinual learning Benchmark

### v1

@@ 3.4 Evaluation Metrics
We evaluate continual learning (CL) models using tailored metrics under a standard CL framework: (1) Average performance metrics: we adopt the final average performance $P$ to test the ability to adapt to sequential video-language tasks. $P$ varies across different tasks, such as MQ, NLQ and VQ. The metric at the $i$-th task is defined as $P_{i}=\frac{1}{i}\sum_{j=1}^{i}p_{i,j}$ where $P$ represents the cumulative performance up to the current task. $p_{i,j}$ is the result evaluated on the j-th task after learned from the i-th task ($j<=i$). In NLQ and MQ tasks, we adopt average recall@k (IoU=m) [14] as the metric, where we select $k=\{1,5\}$ and $m=\{0.3,0.5\}$. This metric presents the percentage of query sentences that appear in the top-k predictions with a tIoU larger than the threshold $m$. For the VQ task, we leverage temporal AP (tAP) [14] to measure the distance between the predictions and ground-truth localizations. (2) Memory stability metrics: we leverage Backward Forgetting (BwF) to calculate the influence caused by the learned task i to the performance of previous tasks. BwF is quantified as $BwF_{i}=\frac{1}{i-1}\sum_{j=1}^{i-1}(p_{j,j}-p_{i,j})$.
@@ 4 Evaluation and Discussion
We evaluate the existing continual learning methods on our newly curated benchmark ViLCo-Bench and examine new query-incremental challenges with multimodal data and tasks in the continual learning setting. We employ several popular continual learning methods, such as EWC [19], MAS [1], iCaRL [30], and BiC [45]. To ensure a fair comparison, all methods utilize the same architectural backbone. For example, we employ EgoVLP-v2 [27] as the visual encoder and CLIP [29] as the text encoder, starting with identical pre-trained parameters for initialization. We report the upper (and lower bound) of the proposed benchmark for each task. These bounds represent multi-task training settings where all samples are available to the model throughout the training process.
@@ 4.2 Ablation Study
Impacts of various backbones. Inspired by Mixture of experts (MoE) [34, 16] in multimodal learning, we evaluate the impacts of different backbones in moment query, such as InternVideo [42], Slowfast [7] and Omnivore [12]. As shown in Table 5, our basic backbone is EgoVLP-v2. We combine EgoVLP-v2 features with the above features by concatenation. It is observed that the combination of EgoVLP-v2 and InternVideo achieves the best performance. Considering that introducing more features increases the difficulty of reducing the gap between video and text representations, more experts are not always better, especially in the continual learning field.
@@
Influence of different task orders. We adopt the same task-splitting strategy with a fixed order. For example, in a moments query, the sub-tasks always involve fixed subsets of action categories. As shown in Table 4.2, we compare different task-splitting strategies by randomly switching orders of tasks. The results demonstrate that replay-based methods are more sensitive to the order of tasks. Due to imbalanced data distribution in different tasks, replay-based methods are easier to overfit to rare categories or queries if the data occurs in the first tasks.
Table 6: Comparing the impact of EM (episodic memory) and SSL (self-supervised learning) modules.
@@
Table 7: Influence of random task order.
@@
EWC 24.2 12.51
iCaRL 4.6 23.66
ViLCo 2.9 26.24
EWC 15.86 18.97
iCaRL 6.21 20.30
ViLCo 1.9 23.02
@@ 5 Discussion and Limitations
Limitation. The current ViLCo-Bench leverages fixed backbones and task-specific heads, optimizing performance for predefined tasks within the CL spectrum. However, there lacks a unified framework that could integrate arbitrary tasks in CL. For instance, tasks such as Visual Question Answering (VQA) [25] and video captioning [38] demand distinct objectives and frameworks that are not inherently supported by ViLCo-Bench’s current structure. Besides architectural limitations, another significant challenge in the current benchmark is the heavy reliance on annotated data streams. Expanding ViLCo-Bench to more modality domains (e.g., audio, images, and time series) exacerbates the requirement for extensive annotations, posing a substantial challenge in real-world applications. Additionally, although data is collected from 9 countries with 74 locations worldwide, half of the videos are from the US, which may introduce societal biases and result in underperformance when applied to other cultural or linguistic groups.
Future works. To further extend our benchmark, we would introduce more sophisticated natural language queries and video-language tasks. We will also include more uni-modal data to learn informative representations, such as depth. Moreover, as we focus on egocentric video and language inputs, ViLCo-Bench would be expanded to related applications, such as household robots. However, the influence of introducing different perspectives (e.g., YouTube videos with a third-person perspective) remains relatively unexplored.
@@ Appendix A Related Works on Continual Learning
Continual learning setups described in the literature are mainly divided into three categories: 1) Task-Incremental (Task-IL), 2) Class-incremental (Class-IL), and 3) domain-incremental (Domain-IL). In Task-IL, the model is trained on different tasks with a similar distribution of inputs [15], while in class incremental, the input distribution changes over time [20, 26]. In domain-incremental, on the other hand, the output distribution from one task to the other task remains the same while the input distribution changes. As the content is provided incrementally in all these scenarios, continual learning is also referred to as incremental learning or lifelong learning in much of the literature, without a strict distinction. This continual/incremental exposure of the model to new data with new distributions leads to the major challenge which is known as catastrophic forgetting, where adaptation to the new data generally results in reducing the ability of the model to remember the old patterns. In fact continual learning is a trade-off between two challenges: ”memory stability“ and ”learning plasticity“. One naive solution to this is to repeat the training process for all the current and previous data, however, this is not possible due to limitations in resources, or missing access to previous data because of privacy. This emphasizes the need for continual and online learning methods. Existing CL approaches to these challenges can be divided into: Regularization-based approach: adding regularization terms with reference to the old model to penalize abrupt changes in the most relevant parameters learned from previous tasks such as Elastic Weight consolidation (EWC) [19]. Replay(Rehearsal)-based: rehearsing and recovering the previous distributions [33, 40]. Architecture-based: customizing architecture to accommodate new tasks by freezing some parameters or sub-layer/networks of the model (WSN[18] or extending the architecture(iCaRL[icarl], BiC). Distilliation-based: Retaining already inference knowledge from previous rounds of training by storing the models or logits and reusing them during new training rounds [4, 20]. We refer to [41] provides a comprehensive study and review of existing continual learning approaches.

### v2

@@ 3.4 Evaluation Metrics
We evaluate continual learning (CL) models using tailored metrics under a standard CL framework following the suggestions by [16, 30]: (1) Average performance metrics: we compare the average performance $P$ to test the ability of the model to adapt to a sequence of video-language tasks. $P$ varies across different tasks, such as MQ, NLQ and VQ. The metric at the $i$-th task is defined as $P_{i}=\frac{1}{i}\sum_{j=1}^{i}p_{i,j}$ where $P$ represents the cumulative performance up to the current task. $p_{i,j}$ is the performance metric evaluated on the i-th task after the model was trained on the previous $j$ tasks ($j<=i$). In NLQ and MQ tasks, we adopt average recall@k (IoU=m) as the performance metric, where we select top $k=\{1,5\}$. This metric presents the percentage of query sentences that appear in the top-k predictions with IoU larger than the threshold $m=\{0.3,0.5\}$. For the VQ task, we leverage temporal AP (tAP) as the performance metric which measures the distance between the predictions and ground-truth localizations. Again we calculate the average of tAP over the previous tasks. (2) Memory stability metrics: Following [46], we also consider Backward Forgetting (BwF) to evaluate the performance of CL models. BwF measures the influence caused by learning task $i$ on the performance of the model in remembering previous tasks. $BwF_{i}$ is calculated as follows: $BwF_{i}=\frac{1}{i-1}\sum_{j=1}^{i-1}(p_{j,j}-p_{i,j})$. in other words, $p_{i,j}$ represents the perforemance of the task $j$ after learning the new task $i$. We report total $BwF=BWF_{N}$ where $N$ is the total number of tasks.
@@ 4 Evaluation and Discussion
We evaluate the existing continual learning methods on our newly curated benchmark ViLCo-BenchẆe employ several popular continual learning methods, such as EWC [22], MAS [1], iCaRL [35], and BiC [51]. To ensure a fair comparison, all methods utilize the same architectural backbone. For example, we employ EgoVLP-v2 [30] as the visual encoder and CLIP [32] as the text encoder, starting with identical pre-trained parameters for initialization. We report the upper (and lower bound) of the proposed benchmark for each task. These bounds represent multi-task training settings where all samples are available to the model throughout the training process.
@@ 4.2 Ablation Study
Impacts of various backbones. Table 5 compares various visual backbones as the vision incoder including EgoVLP-v2, Timersfomer[3] and X3D[8] in extracting effective visual features. Inspired by Mixture of experts (MoE) [39, 19] in multimodal learning, we also evaluated the impacts of different backbones in moment query, such as InternVideo [48], Slowfast [9] and Omnivore [14]. As shown in Table 6, our basic backbone is EgoVLP-v2. We combine EgoVLP-v2 features with the above features by concatenation. It is observed that the combination of EgoVLP-v2 and InternVideo achieves the best performance. Considering that introducing more features increases the difficulty of reducing the gap between video and text representations, more experts are not always better, especially in the continual learning field.
Table 5: The impact of various visual features
[…]
IoU=0.3 IoU=0.5 IoU=0.3 IoU=0.5
Timersformer [3] 2.4 30.80 22.82 51.93 40.64
X3D [8] 1.4 31.50 23.01 48.09 36.59
EgoVLP-v2 [30] 2.9 33.58 26.24 53.75 42.30
@@
Impact of task orders. As shown in Table 4.2, we compare different task-splitting strategies by randomly switching orders of tasks. The results demonstrate that replay-based methods are more sensitive to the order of tasks. Due to imbalanced data distribution in different tasks, replay-based methods are easier to overfit to rare categories or queries if the data occurs in the first tasks. We have provided a more detailed analysis across 6 variations in the order of the tasks in Appendix E
Table 7: Comparing EM (episodic memory) and SSL (self-supervised learning) modules.
@@
Table 8: The impact of the order of tasks.
@@
EWC 1 24.2 12.51
iCaRL 1 4.6 23.66
ViLCo 1 2.9 26.24
EWC 2 15.86 18.97
iCaRL 2 6.21 20.30
ViLCo 2 1.9 23.02
@@ 5 Discussion and Limitations
Limitation. The current ViLCo architecture leverages fixed backbones and task-specific heads, optimized for predefined CL tasks. However, it lacks a unified framework that could integrate arbitrary tasks in CL such as Visual Question Answering (VQA) [28] and video captioning [44] which require distinct objectives and frameworks. Although our proposed model architecture, ViLCo, does not currently include considerations for those tasks, our curated dataset, ViLCo-Bench, remains highly suitable for a wide range of video-language tasks. Besides architectural limitations, another significant challenge in the current benchmark is the heavy reliance on annotated data streams. Expanding ViLCo-Bench to more modality domains (e.g., audio, images, and time series) exacerbates the requirement for extensive annotations, posing a substantial challenge in real-world applications. Additionally, although data is collected from 9 countries with 74 locations worldwide, half of the videos are from the US, which may introduce societal biases and result in underperformance when applied to other cultural or linguistic groups.
Future works. To further extend our benchmark, we would introduce more sophisticated natural language queries and video-language tasks as mentioned above. We will also include more uni-modal data to learn informative representations, such as depth. Moreover, as we focus on egocentric video and language inputs, ViLCo-Bench would be expanded to related applications, such as household robots. However, the influence of introducing different perspectives (e.g., YouTube videos with a third-person perspective) remains relatively unexplored.
@@
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
    2. (b)
      Did you describe the limitations of your work? [Yes] See Section 5.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes] See Section 5.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] Our code and instructions can be found at https://github.com/cruiseresearchgroup/ViLCo. (mentioned on Page 1)
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] Data splits could be found in Section 3.2. Hyperparameters and related details can be seen in Appendix 9
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes] Yes, for MQ tasks (see Table  2), where we run experiments with three different random seeds. For NLQ and VQ tasks, we conduct experiments with a fixed random seed due to training time cost.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See Supplementary material.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes]
    2. (b)
      Did you mention the license of the assets? [Yes] Our benchmark license is shown in the provided GitHub link (mentioned in Page 1). The license of Ego4d (not distributed) is mentioned in the supplementary material.
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes] We provide the download link for new assets in the provided project link (mentioned in Page 1).
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [N/A]
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [N/A]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
@@ Appendix A Related Works on Continual Learning
Continual learning setups described in the literature are mainly divided into three categories: 1) Task-Incremental (Task-IL), 2) Class-incremental (Class-IL), and 3) domain-incremental (Domain-IL). In Task-IL, the model is trained on different tasks with a similar distribution of inputs [18], while in class incremental, the input distribution changes over time [23, 29]. In domain-incremental, on the other hand, the output distribution from one task to the other task remains the same while the input distribution changes. As the content is provided incrementally in all these scenarios, continual learning is also referred to as incremental learning or lifelong learning in much of the literature, without a strict distinction. This continual/incremental exposure of the model to new data with new distributions leads to the major challenge which is known as catastrophic forgetting, where adaptation to the new data generally results in reducing the ability of the model to remember the old patterns. In fact continual learning is a trade-off between two challenges: ”memory stability“ and ”learning plasticity“. One naive solution to this is to repeat the training process for all the current and previous data, however, this is not possible due to limitations in resources, or missing access to previous data because of privacy. This emphasizes the need for continual and online learning methods. Existing CL approaches to these challenges can be divided into: Regularization-based approach: adding regularization terms with reference to the old model to penalize abrupt changes in the most relevant parameters learned from previous tasks such as Elastic Weight consolidation (EWC) [22]. Replay(Rehearsal)-based: rehearsing and recovering the previous distributions [38, 46]. Architecture-based: customizing architecture to accommodate new tasks by freezing some parameters or sub-layer/networks of the model (WSN[21] or extending the architecture(iCaRL[35], BiC[51]). Distilliation-based: Retaining already inference knowledge from previous rounds of training by storing the models or logits and reusing them during new training rounds [5, 23]. We refer to [47] provides a comprehensive study and review of existing continual learning approaches.
@@ Appendix D Training Details
Appendix E Further Analysis on the Impact of the Order of Tasks.
From Figure 5, we observe varying performance based on the order of exposure to different data/sub-tasks. It is important to note that each Task $i$ (ranging from 1 to 5) incorporates different action categories and sample sizes. Sub-tasks with limited data tend to exhibit long-tail distributions, which can cause models to overfit if such tasks are introduced early, leading to performance degradation in subsequent tasks. This phenomenon explains why ViLCo exhibits a higher Avg. R@1 with increased BWF, highlighting that task ordering is a significant source of variance in continual learning.
Figure 5: Impact of the order of tasks in Moment Query scenario: BwF (left) and Average Recall (Avg. R@1) (right)